40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1001     Entry time: Tue Jun 26 16:13:30 2012
Author: Frank 
Type: DailyProgress 
Category: PD 
Subject: power still drifting 

laser power is still drifting with the ISS turned on, but much less. Primary suspects are

  1. sensing PD: is a Silicon PD and so the change in Responsivity caused by thermal fluctuations could be the reason. Sensitivity is typically 1.5%/K @1064nm for Si.
  2. Input offset of SR560

Replaced the PD with a InGaAs diode and re-started a measurement. Will also add sampling of the laser power in the VI. I can't easily switch to the lock-in amplifier as it does not have a GPIB interface to read it.

ELOG V3.1.3-